1.     Explain why, if Cindy equalized the marginal utility per hour of golf and tennis, she would not maximize her utility.

2.     Cindy’s tennis club raises its price of an hour of tennis to $10. The price of golf remains at $10 an hour and Cindy continues to spend $70 on tennis and golf. List the combination of hours spent playing golf and tennis that Cindy can now afford.

3.     Cindy loses her math tutoring job and the amount she has to spend on golf and tennis falls to $35 a month. How does Cindy’s demand for golf change? For Cindy, is golf a normal good or an inferior good? Is tennis a normal good or an infinite good?

 



1
Expert's answer
2022-03-25T15:09:00-0400

Solution

Income= $70/hour

With a value of 10

Maximum value"\frac{70}{10}=7"


Maximum value is 7 hours of golf and no tennis

Price of two goods and the given income is used to Check whether disposable bundle is affordable

"70=xpx+xpx"

"70=T\u00d710+G\u00d710"

"7=T+G"

Therefore the budget suggests it is affordable for the time spent on tennis and golf


The units combinations that Cindy can afford are as follows

0 hours of Tennis and 7 hours of golf

1 hour of tennis and 6 hours of

golf

2 hours of tennis and 5 hours of golf

3 hours of tennis and 4 hours of golf

4 hours of tennis and 3 hours of golf

5hours of tennis and 2 hours of golf

6 hours of tennis and 1 hour of golf

7 hours of tennis and 0 hours of golf


The combination of the budget line is

"( 10\u00d73+10\u00d74=70)"





Cross elasticity

"\frac{∆Qg}{∆pt}×\frac{Pt}{Qg}"


"\frac{-1}{5}×\frac{5}{5}=0.2"


Golf and tennis are compliments, when the demand for golf falls if the price of tennis increases.